OpenStudy (anonymous):

how would you solve (x+1)^3?

OpenStudy (anonymous):

This is an expression. It's not equal to anything

OpenStudy (anonymous):

well, the entire problem is\[r(x+1)=(x+1)^{3}+(x+1)+1\]

OpenStudy (anonymous):

what is r?

OpenStudy (anonymous):

r(x) = x^3 + x + 1

OpenStudy (anonymous):

The problem is r (x+1) - r(x^2)

OpenStudy (anonymous):

(x+1)^3 =x^3 +1 +3x^2 +3x

OpenStudy (anonymous):

this is an algebraic identity i hope nw u cn solve ur xpression by substituting value of (x+1)^3

OpenStudy (anonymous):

well @BrianWest do you no the formula (a+b)^3? well u cud use dat identity..